# DyVal 2: Dynamic Evaluation of Large Language Models by Meta Probing Agents

### v1

@@ Abstract
Evaluation of large language models (LLMs) has raised great concerns in the community due to the issue of data contamination. Existing work designed evaluation protocols using well-defined algorithms for specific tasks, which cannot be easily extended to diverse scenarios. Moreover, current evaluation benchmarks can only provide the overall benchmark results and cannot support a fine-grained and multifaceted analysis of LLMs’ abilities. In this paper, we propose meta probing agents (MPA), a general dynamic evaluation protocol inspired by psychometrics to evaluate LLMs. MPA is the key component of DyVal 2, which naturally extends the previous DyVal (Zhu et al., 2023a). MPA designs the probing and judging agents to automatically transform an original evaluation problem into a new one following psychometric theory on three basic cognitive abilities: language understanding, problem solving, and domain knowledge. These basic abilities are also dynamically configurable, allowing multifaceted analysis. We conducted extensive evaluations using MPA and found that most LLMs achieve poorer performance, indicating room for improvement. Our multifaceted analysis demonstrated the strong correlation between the basic abilities and an implicit Matthew effect on model size, i.e., larger models possess stronger correlations of the abilities. MPA can also be used as a data augmentation approach to enhance LLMs.
@@ 1 Introduction
Intelligence evaluation has never been as important as today due to the contradiction between unprecedented performance and underexplored interpretability of large language models (LLMs) (OpenAI, 2023b; GeminiTeam, 2023). To facilitate a better understanding of the abilities of LLMs, evaluation was carried out by collecting data from various domains (Liang et al., 2023; bench authors, 2023), benchmarking specific tasks (Hendrycks et al., 2021; Chen et al., 2021; Cobbe et al., 2021; Clark et al., 2018), and evaluating in extreme scenarios (Zhu et al., 2023b; Wang et al., 2023; Yang et al., 2022).
@@ 2 Related Work
LLMs Evaluation & Data Contamination. Various benchmarks have been introduced to assess LLMs (Hendrycks et al., 2021; Li et al., 2023a; Zhong et al., 2023; HuggingFace, 2023; Chang et al., 2023), including: (1) Human-centric evaluation, typified by AdaVision (Gao et al., 2022) and AdaTest (Ribeiro and Lundberg, 2022) that emphasize human-driven feedback. (2) Crowd-sourcing, e.g., DynaBench (Kiela et al., 2021) and DynaBoard (Ma et al., 2021), which prioritizes diverse and comprehensive evaluations through crowd-sourced tests. (3) More challenging tasks such as HELM (Liang et al., 2023), DeepTest (Tian et al., 2018) and CheckList (Ribeiro et al., 2020), create custom tests, while platforms such as Big-Bench (bench authors, 2023) aim to challenge LLMs with specialized tasks.
@@
Recent research has highlighted the significant issue of data contamination in LLMs (Lovin, 2023; Bender et al., 2021; Kocoń et al., 2023; Li, 2023; Zhou et al., 2023). In particular, the reports of GPT-4 (OpenAI, 2023b), LLama (Touvron et al., 2023), and Skywork LLM (Wei et al., 2023) have acknowledged this phenomenon. Several researchers (Golchin and Surdeanu, 2023a, b; Oren et al., 2023; Yang et al., 2023; Oren et al., 2023) developed methods to detect data contamination. Zhu et al. (2023a); Lei et al. (2023); Fan et al. (2023) introduced dynamic evaluation strategies via different algorithms to reduce data contamination. MPA significantly differs from them in the sample generation mechanism and the support for multifaceted analysis.
LLMs as Autonomous Agents. The adoption of LLMs as autonomous agents for task completion has recently gained popularity, such as AutoGPT (Significant-Gravitas, 2023) and MetaGPT (Hong et al., 2023), which have advanced our understanding of collaborative and planning abilities in LLMs. Another growing trend is the use of LLMs as judges (Dubois et al., 2023; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2023; Wang et al., 2024), where LLMs assess the output of other LLMs. Furthermore, there is a burgeoning interest in leveraging LLMs to enhance training datasets (Wang et al., 2022; Yu et al., 2023; Yuan et al., 2024; Liu and Yao, 2024) and aligning LLMs’ outputs with specific goals or values (Burns et al., 2023). Our work differs from them in designing psychometrically inspired principles for sample generation and support for multifaceted analysis.
@@ 3.1 Overview
In this work, we take inspiration from psychometrics (Raykov and Marcoulides, 2011; Burnell et al., 2023) to generate evaluation samples using LLMs as agents. Specifically, instead of relying on specific rules like DyVal (Zhu et al., 2023a), we employ LLMs as agents to automatically generate new problems based on the given evaluation samples. This agent-based evaluation design can potentially fit most tasks. More importantly, psychometric theory categorizes cognitive abilities into three basic ones: language understanding to comprehend and generate texts, problem-solving to deal with complex problems, and domain knowledge. Therefore, as shown in Figure 2(a), we follow these criterion to design principles that not only change the problems but also assist in multifaceted analysis.
@@ 3.4 Psychometric principles
Psychometric principles guide how we probe the understanding of a question. Inspired by psychometric theory (Raykov and Marcoulides, 2011), we aim to evaluate three basic abilities of LLMs: language understanding, problem solving, and domain knowledge. We have identified five key principles that correspond to these categories, as shown in Table 1. These principles incorporate both a probing agent and a judge agent, as previously discussed, with their functions differing according to the specific principle applied.
@@ 4.3 Effect of Different Probing Principles
The results on MMLU and ARC-C datasets are detailed in Figure 4 and Section C.2. It can be observed that, principle $1,3$ and $4$ are most effective principles. While principle $3$, which randomly permute choices, are less effective. For GPT-4-Turbo and GPT-3.5-Turbo on ARC-C dataset, it can intriguingly increase the performance.
@@ 4.4 Ablation Study on Prompt Engineering
[…]
@@
[…]
@@ 5.4 Error Analysis
  • Domain knowledge. We investigated the distribution of topic error among $57$ tasks of MMLU in Section C.4. Notably, the professional law domain has the highest error rate, followed by moral scenarios and professional psychology, suggesting challenges predominantly in professional and ethical tasks.
@@ 6 MPA as Data Augmentation for LLMs
Although the main purpose of MPA is to evaluate LLMs, its generated samples can also be used as augmented data for fine-tuning. In this section, we conducted a pilot study using the OpenAI API to fine-tune GPT-3.5-Turbo on the data generated by MPA. Specifically, we used samples from the training split of MMLU and ARC-C, which are fed to MPA for data generation. We then evaluated the fine-tuned models on the original test split and our probing benchmarks. The results in Figure 5 show that the data generated by MPA can improve the performance of LLMs, with an average of $2\%$ improvements on both MMLU and ARC-C. The fine-tuning results demonstrated that MPA is not only an evaluation protocol, but a general data augmentation approach to improve the performance of LLMs, creating a huge advantage for training stronger LLMs in the future. The improved results also demonstrate the correctness of the generated data, indicating that our MPA is effective.
@@ C.1 Standard Deviation of Main Results
Table 5: The standard deviation of efficient of the main results.

### v2

# Dynamic Evaluation of Large Language Models by Meta Probing Agents

@@ Abstract
Evaluation of large language models (LLMs) has raised great concerns in the community due to the issue of data contamination. Existing work designed evaluation protocols using well-defined algorithms for specific tasks, which cannot be easily extended to diverse scenarios. Moreover, current evaluation benchmarks can only provide the overall benchmark results and cannot support a fine-grained and multifaceted analysis of LLMs’ abilities. In this paper, we propose meta probing agents (MPA), a general dynamic evaluation protocol inspired by psychometrics to evaluate LLMs. MPA designs the probing and judging agents to automatically transform an original evaluation problem into a new one following psychometric theory on three basic cognitive abilities: language understanding, problem solving, and domain knowledge. These basic abilities are also dynamically configurable, allowing multifaceted analysis. We conducted extensive evaluations using MPA and found that most LLMs achieve poorer performance, indicating room for improvement. Our multifaceted analysis demonstrated the strong correlation between the basic abilities and an implicit Matthew effect on model size, i.e., larger models possess stronger correlations of the abilities. MPA can also be used as a data augmentation approach to enhance LLMs. Code is available at: https://github.com/microsoft/promptbench.
LLMs, Evaluation
[…]
@@ 1 Introduction
Intelligence evaluation has never been as important as today due to the contradiction between unprecedented performance and underexplored interpretability of large language models (LLMs) (OpenAI, 2023b; GeminiTeam, 2023). To facilitate a better understanding of the strength and weakness of LLMs, evaluation was carried out by collecting data from various domains (Liang et al., 2023; bench authors, 2023), benchmarking specific tasks (Hendrycks et al., 2021; Chen et al., 2021; Cobbe et al., 2021; Clark et al., 2018), and evaluating in extreme scenarios (Zhu et al., 2023b; Wang et al., 2023; Yang et al., 2022).
@@ 2 Related Work
LLMs Evaluation & Data Contamination. Various benchmarks have been introduced to assess LLMs (Hendrycks et al., 2021; Li et al., 2023a; Zhong et al., 2023; HuggingFace, 2023; Chang et al., 2023), including: (1) Human-centric evaluation, typified by AdaVision (Gao et al., 2022) and AdaTest (Ribeiro & Lundberg, 2022) that emphasize human-driven feedback. (2) Crowd-sourcing, e.g., DynaBench (Kiela et al., 2021) and DynaBoard (Ma et al., 2021), which prioritizes diverse and comprehensive evaluations through crowd-sourced tests. (3) More challenging tasks such as HELM (Liang et al., 2023), DeepTest (Tian et al., 2018) and CheckList (Ribeiro et al., 2020), create custom tests, while platforms such as Big-Bench (bench authors, 2023) aim to challenge LLMs with specialized tasks.
@@
Recent research has highlighted the significant issue of data contamination in LLMs (Lovin, 2023; Bender et al., 2021; Kocoń et al., 2023; Li, 2023; Zhou et al., 2023). In particular, the reports of GPT-4 (OpenAI, 2023b), LLama (Touvron et al., 2023), and Skywork LLM (Wei et al., 2023) have acknowledged this phenomenon. Several researchers (Golchin & Surdeanu, 2023a, b; Oren et al., 2023; Yang et al., 2023; Oren et al., 2023) developed methods to detect data contamination. Zhu et al. (2023a); Lei et al. (2023); Fan et al. (2023) introduced dynamic evaluation strategies via different algorithms to reduce data contamination. MPA significantly differs from them in the sample generation mechanism and the support for multifaceted analysis.
LLMs as Autonomous Agents. The adoption of LLMs as autonomous agents for task completion has recently gained popularity, such as AutoGPT (Significant-Gravitas, 2023) and MetaGPT (Hong et al., 2023), which have advanced our understanding of collaborative and planning abilities in LLMs. Another growing trend is the use of LLMs as judges (Dubois et al., 2023; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2023; Wang et al., 2024), where LLMs assess the output of other LLMs. Furthermore, there is a burgeoning interest in leveraging LLMs to enhance training datasets (Wang et al., 2022; Yu et al., 2023; Yuan et al., 2024; Liu & Yao, 2024) and aligning LLMs’ outputs with specific goals or values (Burns et al., 2023). Our work differs from them in designing psychometrically inspired principles for sample generation and support for multifaceted analysis.
@@ 3.1 Overview
In this work, we take inspiration from psychometrics (Raykov & Marcoulides, 2011; Burnell et al., 2023) to generate evaluation samples using LLMs as agents. Specifically, instead of relying on specific rules like DyVal (Zhu et al., 2023a), we employ LLMs as agents to automatically generate new problems based on the given evaluation samples. This agent-based evaluation design can potentially fit most tasks. More importantly, psychometric theory categorizes cognitive abilities into three basic ones: language understanding to comprehend and generate texts, problem-solving to deal with complex problems, and domain knowledge. Therefore, as shown in Figure 2(a), we follow these criterion to design principles that not only change the problems but also assist in multifaceted analysis.
@@ 3.3 Judge Agent
Human Verification
For each ability (language understanding, problem solving, and domain knowledge), we randomly selected $500$ samples from the MMLU dataset and $100$ samples from the ARC-c dataset, totaling $1,800$ questions. $30$ human experts (with bachelor or higher degree) are divided into 3 groups, each with 10 person. They were asked to judge the following two questions: (1) whether the original and paraphrased questions were equivalent; (2) if the answers to the probing questions remained correct. The evaluation required a simple ‘Yes’ or ‘No’ answer.
The positive results shown in Appendix Table 9 from our human evaluation with an overall accuracy rate of $94$% and $97$% for each question, underscoring the effectiveness of our methodology. The following table provides a detailed breakdown of the evaluation outcomes, showcasing the high level of confidence in the equivalence and correctness of our probing questions across the three abilities.
@@ 3.4 Psychometric principles
Psychometric principles guide how we probe the understanding of a question. Inspired by psychometric theory (Raykov & Marcoulides, 2011), we aim to evaluate three basic abilities of LLMs: language understanding, problem solving, and domain knowledge. We have identified five key principles that correspond to these categories, as shown in Table 1. These principles incorporate both a probing agent and a judge agent, as previously discussed, with their functions differing according to the specific principle applied.
@@ 4.1 Experimental Setup
A bitter reality is that currently, only GPT-4 are capable to generate questions and judge the quality of generated questions. We believe as the field progresses, more cheaper models (such as Claude 3 and Gemini) will become capable of fulfilling both probing and judging roles, thereby reducing dependency on any single model’s API and enhancing the scalability.
@@ 4.3 Effect of Different Probing Principles
The results on MMLU and ARC-C datasets are detailed in Figure 4 and Section C.2. It can be observed that, principle $1,2$ and $5$ are the most effective principles. While principle $3$, which randomly permute choices, are less effective. For GPT-4-Turbo and GPT-3.5-Turbo on ARC-C dataset, it can intriguingly increase the performance.
@@ 4.4 Ablation Study on Prompt Engineering
We explored the efficacy of prompt engineering techniques, specifically Chain-of-Thought (CoT) (Wei et al., 2022) and In-Context Learning (ICL) (Brown et al., 2020). For ICL, we selected five examples from the corresponding training set as few-shot examples. We also applied MPA to these five examples, creating a transformed version of few-shot examples. We refer to the use of ICL with the original training examples as ICLo and the use of ICL with the transformed examples as ICLt.
@@
4.5 Albation Study on Data Contamination
We collected $30$ novel reasoning questions from the experts (the same experts in the human evaluation in the general response), and examined the performance of the original questions and the probing questions. The GPT-4’s accuracies of the original questions and the probing questions are both $60$%. The OT/PF ratio is $3$%, which is much lower than those in common public benchmarks, indicating that the newly generated benchmarks are not likely to be memorized compared to the public benchmarks.
@@ 5.4 Error Analysis
  • Domain knowledge. We investigated the distribution of topic error among $57$ tasks of MMLU in Figure 10. Notably, the professional law domain has the highest error rate, followed by moral scenarios and professional psychology, suggesting challenges predominantly in professional and ethical tasks.
@@
Figure 8: The bar chart of top 20 error topics and their corresponding frequencies of GPT-4-Turbo on MMLU dataset.
@@ 6 MPA as Data Augmentation for LLMs
Although the main purpose of MPA is to evaluate LLMs, its generated samples can also be used as augmented data for fine-tuning. In this section, we conducted a pilot study using the OpenAI API to fine-tune GPT-3.5-Turbo on the data generated by MPA. Specifically, we used samples from the training split of MMLU and ARC-C, which are fed to MPA for data generation. We then evaluated the fine-tuned models on the original test split and our probing benchmarks. The fine-tuning data includes two parts, the first part is the probing questions generated by $5$ principles separately, and the second part is the original training set.
The results in Figure 9 show that the data generated by MPA can improve the performance of LLMs, with an average of $2\%$ improvements on both MMLU and ARC-C. The fine-tuning results demonstrated that MPA is not only an evaluation protocol, but a general data augmentation approach to improve the performance of LLMs, creating a huge advantage for training stronger LLMs in the future. The improved results also demonstrate the correctness of the generated data, indicating that our MPA is effective.
@@ C.1 Standard Deviation of Main Results
Table 5: The standard deviation of co-efficient of the main results.
@@ Appendix D Examples generated by MPA
Appendix E Human verification results
Table 9: Results of human verification on MMLU and ARC-Challenge datasets.
Equivalence / Correctness Language Understanding Problem Solving Domain Knowledge Avg
Group 1 0.88/0.95 0.91/0.93 1.00/1.00 0.93/0.96
Group 2 0.92/0.99 0.94/0.98 1.00/0.97 0.95/0.98
Group 3 0.91/0.95 0.92/0.99 1.00/1.00 0.94/0.98
Avg 0.90/0.96 0.92/0.97 1.00/0.99 0.94/0.97